# Identifying treatment response subgroups in observational time-to-event data

### v3

@@ 2.2 Estimating the quantities of interest with neural networks
The previous section discussed the quantities one must estimate —here parameterised by neural networks— to uncover subgroups of treatment effects: the assignment function $\mathbb{P}(Z\mid X)$, the cumulative incidence $\Lambda_{k}$ characterising survival under the two treatment regimes, and the weights $w$. Figure 3 illustrates the overall architecture and the neural networks used to estimate these quantities.
@@
Similar to Jeanselme et al. (2022), a multi-layer perceptron $G$ with final Softmax layer assigns each patient characterised by covariates $x$ its probability of belonging to each subgroup, characterised through a $K$-dimensional vector of probabilities.
@@
Each subgroup $k$ is represented by a vector $l_{k}\in\mathbbm{R}^{L}$ of dimension $L$, a latent parametrisation of the cumulative hazard functions. The vector $l_{k}$ is concatenated with $t$ and used as input to a neural network $M$ with monotonic positive outcomes [^3] To ensure this constraint, we apply a square function on all weights as proposed in Jeanselme et al. (2022; 2023)., with a final SoftPlus layer to model the cumulative hazard functions under both treatment regimes $\Lambda_{k}(t):=(\Lambda_{k}(t\mid A=0),\Lambda_{k}(t\mid A=1))$. We use the following transformation to ensure that no probability is assigned to negative times, a limitation raised concerning previous monotonic neural networks (Shchur et al., 2020):
$\Lambda_{k}(t):=t\cdot M(l_{k},t)$
By modelling these two cumulative intensity functions, we can estimate the Subgroup Average Treatment Effect, $\hat{\tau}_{k}(t)$, in subgroup $k$ as
$\hat{\tau}_{k}(t):=\mathbb{E}[\tau(t,X)\mid Z=k)]=\exp(-\Lambda_{k}(t\mid A=1))-\exp(-\Lambda_{k}(t\mid A=0))$
@@
Under treatment randomisation, as in RCTs, the previous components would accurately model the observed survival outcomes and identify subgroups of treatment effects by maximising the factual likelihood. As previously discussed, to account for the treatment non-randomisation in observational studies, we weight the factual likelihood using the propensity score of a patient estimated through a multi-layer perceptron $W$ with a final sigmoid transformation as
@@ 2.3 Training procedure
First, the network $W$ is trained to predict the binary treatment assignment by minimising the cross-entropy of receiving treatment. Then, training all other components relies on minimising the weighted factual log-likelihood introduced in equation 6. The use of monotonic neural networks results in the efficient and exact computation of the log-likelihood as automatic differentiation of the monotonic neural networks’ outcomes readily provides the instantaneous hazards $\lambda_{k}$ necessary for computing the survival function derivative in the log likelihood (Jeanselme et al., 2022; 2023; Rindt et al., 2022).
@@ 3 Experimental analysis
As counterfactuals are unknown in observational data, we adopt—as is common practice in this research field—a synthetic dataset in which underlying survival distributions and group structure are known. Appendix C accompanies these synthetic results with the analysis of heterogeneity in real-world, adjuvant radiotherapy responses for patients diagnosed with breast cancer. Our code to produce the synthetic dataset, the model, and the reproduction of all experimental results is available on Github [^4] https://github.com/Jeanselme/CausalNeuralSurvivalClustering.
@@ 3.1 Data generation
We generate a population of $3,000$ equally divided into $K=3$ subgroups. We draw 10 covariates from normal distributions with different centres and survival distributions for each treatment regime following Gompertz distributions (Pollard & Valkovics, 1992) parameterised by group membership and individual covariates. Note that this setting relaxes the independence of outcomes and covariates assumed by the proposed model. This choice reflects a setting in which time-to-event distributions under treatment or control regimes are different functions of the covariates. This simulation is more likely to capture the complexity of real-world responses, in contrast to traditional evaluations of subgroup analysis which often assume a linear treatment response, i.e., a shift between treated and untreated distributions. Further, we implement two treatment assignment scenarios: a randomised assignment in which treatment is independent of patient covariates, similar to RCTs (Randomised), and one in which treatment is a function of the patient covariates as in observational studies (Observational). Finally, non-informative censoring times are drawn following a Gompertz distribution. Details of the generative process of our synthetic dataset are deferred to App. B.1.
@@ 3.2 Empirical settings
We compare the proposed approach against the state-of-the-art Cox Mixtures with Heterogeneous Effect (CMHE Nagpal et al. (2022a) [^5] Implemented in the Auton-Survival library (Nagpal et al., 2022b)) which uncovers treatment effect and baseline survival latent groups. This method uses an expectation maximisation framework in which each patient is assigned to a group for which a Cox model is then fitted. The central differences with our proposed approach is that CMHE (i) clusters patients in treatment effects subgroups and survival subgroups, and (ii) assumes a linear treatment response. This separation between survival and treatment response improves the model’s flexibility at the cost of interpretability as the number of groups grows exponentially, and subgroups of treatment effect are independent of survival. Further, the assumption of linear treatment response may hinder the discovery of subgroups with more complex responses. By contrast, our approach identifies subgroups of treatment while considering survival, without constraining the treatment response. We argue that these are key strengths of our method as a group not responding to treatment with low life expectancy would most benefit from alternative treatments, in comparison to a group with the same treatment responses but better survival odds. Considering jointly non-linear treatment effects and survival, therefore, results in identifying more clinically relevant subgroups.
@@
For a fair comparison, we present three alternatives of CMHE: one with fixed $K=3$ survival subgroups, one with $L=3$ treatment effect subgroups, and one with $K=L=2$, which allows for a total of 4 subgroups. Crucially, these methods assume proportional hazards for each subgroup and do not account for the treatment non-randomisation. Additionally, we also compare our model against its unadjusted alternative (CSC Unadjusted), which uses the unweighted factual likelihood ($w_{i}=1$) and should suffer from non-randomisation of treatment assignment.
@@
Finally, we compare against two step-wise approaches in which clustering and treatment effect are modelled separately. First, we use an unsupervised clustering algorithm on the covariates, followed by a non-parametric estimate of the treatment effect as proposed in Nagpal et al. (2022a), referred as Kmeans + TE in the following. Specifically, we use a K-Means (Hartigan & Wong, 1979) to cluster the data and we compute the difference between Kaplan-Meier (Kaplan & Meier, 1958) estimates between control and treated patients stratified by clusters. However, this approach is limited by its implicit assumption of randomisation and its pre-specified number of subgroups, which is not informed by observed outcomes. Second, we use a Virtual Twins approach to estimate individualised treatment effects and then cluster them. As proposed in the literature, we propose to use a survival tree to estimate response under each treatment regime and then use a KMeans on the estimated treatment effects, computed as the difference between survival estimates. Note that this methodology also assumes treatment randomisation, and the clustering is uninformed by the covariates. For details on training and hyperparameter optimisation, we refer to App. B.2.
@@ 3.3 Treatment effect recovery
In this section, we present the results under the previously described data generation process. App. B.4 presents alternative datasets which demonstrate the flexibility of the proposed methodology.
Figure 4: Averaged negative log-likelihood across 5-fold cross-validation given the number of subgroups $K$ under the "Observational" treatment assignment with the shaded area representing 95% CI. The log-likelihood presents an elbow around the underlying number of subgroups.
@@
For all methodologies, we need to choose the number of subgroups a priori. An important question is, therefore, whether we can identify the true number of subgroups with our model. Figure 4 presents the average NLL obtained by cross-validation for models with different numbers of subgroups $K$. The dotted lines represent the elbow heuristic (Thorndike, 1953), which identifies a change point in the explained variability of a clustering strategy, here considering the log-likelihood. Using this heuristic, the optimal choice for $K$ is $3$, which agrees with the underlying generative process. App. B.3 explores the sensitivity to this choice, demonstrating the method’s robustness. This data-driven choice of the number of subgroups $K$ is a crucial strength of our method, compared to classical two-stage analyses which separate clustering from treatment effect estimates. In these methods, survival outcomes cannot directly guide the choice of $K$.
@@
Next, we present quantitative results of our method against the benchmarks methods. Table 1 presents the performance of the different methodologies under the two studied scenarios. Recall that $K$ denotes the number of treatment response subgroups, while in CMHE, the additional parameter $L$ describes the number of survival distributions.
@@
Model Rand-Index $\text{IAE}_{k}(t_{max})$ $k=1$ $k=2$ $k=3$ Randomised CSC 0.798 (0.054) 0.042 (0.012) 0.032 (0.007) 0.015 (0.006) CSC Unadjusted 0.804 (0.052) 0.032 (0.009) 0.027 (0.007) 0.018 (0.010) CMHE ($K=3$) 0.392 (0.034) 0.164 (0.009) 0.077 (0.004) 0.070 (0.005) CMHE ($L=3$) 0.258 (0.155) - 0.088 (0.007) - CMHE ($K=L$) 0.392 (0.054) 0.338 (0.262) - - KMeans + TE 0.000 (0.002) 0.210 (0.011) 0.091 (0.007) 0.142 (0.016) Virtual Twins 0.578 (0.081) 0.022 (0.006) 0.032 (0.011) 0.025 (0.010) Observational CSC 0.812 (0.038) 0.037 (0.013) 0.034 (0.009) 0.022 (0.005) CSC Unadjusted 0.727 (0.084) 0.048 (0.019) 0.025 (0.011) 0.025 (0.009) CMHE ($K=3$) 0.385 (0.022) 0.169 (0.012) 0.078 (0.005) 0.075 (0.008) CMHE ($L=3$) 0.190 (0.127) 0.192 (0.010) - 0.140 (0.009) CMHE ($K=L$) 0.454 (0.068) 0.210 (0.013) 0.095 (0.020) 0.188 (0.014) KMeans + TE 0.001 (0.004) 0.192 (0.016) 0.090 (0.019) 0.147 (0.016) Virtual Twins 0.438 (0.139) 0.050 (0.039) 0.034 (0.033) 0.051 (0.033)
@@
Table 1: 5-fold cross-validated performance averaged (with standard deviation in parenthesis) under the Randomised and Observational treatment simulation. Best performance per column and simulation scenario are marked in bold, second best in italic. ’-’ describes when the methodology diverges. Our proposed CSC method and its unadjusted variant best recovers the underlying treatment responses, with the adjusted approach presenting the best performance in the observational setting.
@@
CSC outperforms all CMHE alternative, the current state-of-the-art method. CMHE’s parametrisation and assumptions explain this difference. Critically, CMHE assumes (i) proportional hazards and (ii) a linear impact of treatment on log-hazards. Neither of these two assumptions is likely to hold in real-world settings as mimicked by our synthetic data. In contrast, our approach does not constrain the treatment effect due to its flexible modelling of the survival function under both treatment regimes. Increasing the number of subgroups, as shown in $K=L$ improves the performance of CMHE in terms of clustering quality (Rand-Index) and the recovery of the underlying treatment effect (lower IAE), but still is inferior compared to our proposed methods. These experiments highlight the advantage of the proposed CSC method in uncovering subgroups of treatment responses due to the flexibility in modelling complex survival distributions under both treatment regimes without proportionality assumption.
@@
CSC presents the best performance in identifying the underlying subgroup. Our proposed CSC method outperforms all approaches on the Rand-Index. In particular, the two-step approaches do not identify the underlying subgroups when covariates and outcomes subgroups are aligned. KMeans + TE identify subgroups that are independent of the outcome of interest, as shown by the Rand-Index. Consequently, the identified clusters differ in their treatment responses as shown by the large IAE. The Virtual Twins approach presents a better capacity to identify the clusters of interest with better Rand-Index and IAE. However, the two-step approach hurts subgroup identification due to the disconnect between the clustering and input covariates. This disconnect leads to a significantly lower Rand-Index in comparison to CSC despite accurate treatment effect estimates.
@@
Treatment assignment correction improves subgroup identification in observational settings. The Observational simulation demonstrates the importance of correcting the likelihood under treatment non-randomisation. The two CSC alternatives present comparable performance in the randomised setting, as theoretically expected, due to a constant $w_{i}$ in this context. Note that estimating propensity slightly reduces performance due to error introduced, reducing performance for the corrected alternative. In a randomised setting, CSC Unadjusted should be preferred. However, the superiority of CSC over its unadjusted alternative is evident in the observational setting, where CSC better recovers the different groups as shown by the Rand-Index and subgroup treatment effects. Finally, Appendices B.4.1 to B.4.5 further validate the robustness of our method in settings with unequally distributed cluster sizes, population size, increased number of underlying clusters, different treatment rates and covariate structures.
@@
Using the Surveillance, Epidemiology, and End Results program [^7] Available at https://seer.cancer.gov/(Seer), App. C investigates a real-world setting: we analyse which subgroups of breast cancer patients most benefit from adjuvant radiotherapy, an active area of research. Our analysis highlights a subgroup characterised by a higher number of distant lymph nodes that benefit from treatment. Despite the inherent limitations of the dataset, our exploratory method identifies a subgroup that alternative methods miss, potentially informing future clinical trials to validate or challenge this hypothesis. Critically, this demonstrates that our work offers a practical methodology for exploratory subgroup analysis of medical data to identify treatment effect heterogeneity in observational studies.
4 Related work
While survival subgrouping has been proposed through mixtures of distributions (Nagpal et al., 2021a; b; Jeanselme et al., 2022) to identify different phenotypes of patients, the machine learning literature on phenotyping treatment effects with time-to-event outcomes remains sparse. The current literature focuses on estimating population or individual treatment effects (Curth et al., 2021; Shalit et al., 2017; Johansson et al., 2016; Zhang et al., 2020). While these models help to understand the average population-wide response to treatment and aim to estimate individualised treatment effects, they do not provide an understanding of the patient groups that may benefit more or be harmed more by treatment. Identifying such groups aligns with, and is thus more useful for, drafting medical guidelines that direct treatment to those subgroups most likely to benefit from it.
Tree-based approaches were proposed to address the limitations of step-wise approaches by jointly discovering subgroups and modelling the treatment effect. Instead of traditional splits on the observed outcomes, these causal trees aim to discover homogeneous splits regarding covariates and treatment effects. Su et al. (2009) introduce a recursive population splitting based on the average difference in treatment effect between splits. Athey & Imbens (2015; 2016) improve the confidence interval estimation through the honest splitting criterion, which dissociates the splitting from the treatment effect estimation. Wager & Athey (2018) agglomerate these causal trees into causal forests for improved ITE estimation. Each obtained split in the decision tree delineates two subgroups of treatment effect Lipkovich et al. (2011); Loh et al. (2015). Alternatively, McFowland III et al. (2018) propose pattern detection and Wang & Rudin (2022), causal rule set learning to uncover these subgroups. However, all these approaches rely on a local optimisation criterion (Lipkovich & Dmitrienko, 2014) and greedy split exploration. Recently, Nagpal et al. (2020) addressed the local optimisation by constraining the treatment response to a linear form in a mixture of Cox models.
Previous approaches uncover subgroups of treatment effect but consider RCTs with binary outcomes, not the observational setting with inherently continuous survival outcomes that our paper explores. At the intersection with survival analysis, Zhang et al. (2017) extend causal trees to survival causal trees, modifying the splitting criterion by measuring the difference in survival estimates between resulting leaves. Similarly, Hu et al. (2021) propose Bayesian additive models and Zhu & Gallego (2020) propose a step-wise approach with propensity weighting to study observational data. Closest to our work, (Jia et al., 2021; Nagpal et al., 2023) propose to uncover subgroups within RCTs with survival outcomes. Jia et al. (2021) propose a mixture of treatment effects characterised by Weibull distributions trained in an expectation-maximisation framework. Similarly, Nagpal et al. (2023) stratify the population into three groups: non-, positive- and negative responders to treatment. An iterative Monte Carlo optimisation is used to uncover these subgroups, characterised by a Cox model with a multiplicative treatment effect. As demonstrated in our work, this step-wise optimisation may be limiting, and the assumption of RCTs renders the model less relevant in observational data.

### v4

@@ 3.2 Estimating the quantities of interest
The previous section discussed the quantities one must estimate —here parametrised by neural networks [^2] When interpretability is a key concern, one can consider a linear model for both subgroup assignment and IPW components, as explored in Appendix B.4.— to uncover subgroups of treatment effects: the assignment function $\mathbb{P}(Z\mid X)$, the distributions characterising survival under the two treatment regimes $\mathbb{P}(T^{\prime}\geq t\mid A,Z=k)$, and the IPW weights $w$. Figure 3 illustrates the overall architecture and the neural networks used to estimate these quantities.
@@
Similar to (Jeanselme et al., 2022), a multi-layer perceptron $G$ with a final Softmax layer assigns each patient characterised by covariates $x$ its probability of belonging to each subgroup, characterised through a $K$-dimensional vector of probabilities.
@@
Each subgroup $k$ is represented by a parameter vector $u_{k}\in\mathbbm{R}^{L}$ of dimension $L$, a latent parametrisation of the survival functions learnt through backpropagation. The vector $u_{k}$ is concatenated with $t$ and used as input to a neural network $M$ with monotonic outcomes [^3] To ensure this constraint, we apply a square function on all weights as proposed in (Jeanselme et al., 2022, 2023). and a final SoftPlus layer to ensure positivity. We train this neural network to model survival through the following transformation — ensuring that no probability is assigned to negative times, a limitation raised concerning previous monotonic neural networks (Shchur et al., 2020):
$\mathbb{P}(T^{\prime}_{A=a}\geq t\mid Z=k):=\exp(-t\cdot M(u_{k},t)[a])$
@@
Under treatment randomisation, as in RCTs, the previous components would accurately model the survival outcomes under both treatment regimes and identify subgroups of treatment effects by maximising the factual likelihood. As previously discussed, to account for the treatment non-randomisation in observational studies, we weight the factual likelihood using the propensity of receiving treatment estimated through a multi-layer perceptron $W$ with a final sigmoid transformation as:
@@
From this estimated probability, we further truncate the propensity score (Austin & Stuart, 2015) to avoid extreme values for the estimated weights, which would result in unstable treatment effect estimates. In this context, the weights $w_{i}$ are defined as:
$\forall i,w_{i}^{\prime}:=$ $a_{i}W(x_{i})+(1-a_{i})(1-W(x_{i}))$
$w_{i}:=$ $\max(0.05,\min(w_{i}^{\prime},0.95))$ (4)
@@ 3.3 Training procedure
First, the model $W$ is trained to predict the binary treatment assignment probability by minimising the cross-entropy of receiving treatment. Then, training all other components relies on maximising the weighted log-likelihood introduced in Equation (2). The use of monotonic neural networks results in the efficient and exact computation of the log-likelihood as a forward path provides the estimated survival and automatic differentiation of the monotonic neural networks’ outcomes readily provides the derivative necessary for computing the log-likelihood (Jeanselme et al., 2022, 2023; Rindt et al., 2022).
@@ 4 Experimental analysis
As counterfactuals are unknown in observational data, we adopt—as is common practice in this research field—a synthetic dataset in which underlying survival distributions and group structure are known. As a real-world case study, Appendix C accompanies these synthetic results with the analysis of heterogeneity in adjuvant radiotherapy responses for patients diagnosed with breast cancer. Our code to produce the synthetic dataset, the model, and all experimental results is available on Github [^4] https://github.com/Jeanselme/CausalNeuralSurvivalClustering.
@@ 4.1 Data generation
We generate a population of $3,000$ equally divided into $K=3$ subgroups. We draw 10 covariates from normal distributions with different centres and survival distributions for each treatment regime following Gompertz distributions (Pollard & Valkovics, 1992) parametrised by group membership and individual covariates. Note that this setting relaxes the independence of outcomes and covariates assumed by the proposed model (Assumption 2). This choice reflects a setting in which survival distributions under treatment or control regimes are different functions of the covariates. This simulation is more likely to capture the complexity of real-world responses, in contrast to traditional evaluations of subgroup analysis, which often assume a linear treatment response. Further, we implement two treatment assignment scenarios: a randomised assignment in which treatment is independent of patient covariates, similar to RCTs (Randomised), and one in which treatment is a function of the patient covariates as in observational studies (Observational). Finally, non-informative censoring times are drawn following a Gompertz distribution. Details of the generative process of our synthetic dataset are deferred to Appendix B.1. Further, Appendix B.5 explores alternative datasets to demonstrate the flexibility of the proposed methodology.
@@ 4.2 Empirical settings
We compare the proposed approach (CSC) against the state-of-the-art Cox Mixtures with Heterogeneous Effect (CMHE (Nagpal et al., 2022a) [^5] Implemented in the Auton-Survival library (Nagpal et al., 2022b)) which uncovers treatment effect and baseline survival latent groups. This method uses an expectation maximisation framework in which each patient is assigned to a group for which a Cox model is then fitted. The central differences with our proposed approach are that CMHE (i) clusters patients in treatment effects subgroups and survival subgroups, and (ii) assumes a linear treatment response. This separation between survival and treatment response improves the model’s flexibility at the cost of interpretability as the number of groups grows exponentially, and subgroups of treatment effect are independent of survival. Further, the assumption of linear treatment response may hinder the discovery of subgroups with more complex responses. By contrast, our approach identifies treatment subgroups while considering survival without constraining the treatment response. We argue that these are key strengths of our method as a group not responding to treatment with low life expectancy would most benefit from alternative treatments, compared to a group with the same treatment responses but with better survival odds. Considering jointly non-linear treatment effects and survival, therefore, results in identifying more clinically relevant subgroups.
@@
For a fair comparison, we present three alternatives of CMHE: one with fixed $K=3$ survival subgroups, one with $L=3$ treatment effect subgroups, and one with $K=L=2$, which allows for a total of 4 subgroups. Crucially, these methods assume proportional hazards for each subgroup and do not account for the treatment non-randomisation.
Additionally, we also compare our model against its unadjusted alternative (CSC Unadjusted), which uses the unweighted factual likelihood ($w_{i}=1$) and should suffer from non-randomisation of treatment assignment.
@@
Finally, we compare against two step-wise approaches in which clustering and treatment effect are modelled separately. First, we use an unsupervised clustering algorithm on the covariates, followed by a non-parametric estimate of the treatment effect as proposed in (Nagpal et al., 2022a), referred to as Kmeans + TE in the following. Specifically, we use a K-Means (Hartigan & Wong, 1979) to cluster the data, and we compute the difference between Kaplan-Meier (Kaplan & Meier, 1958) estimates between control and treated patients stratified by clusters. Second, we use a Virtual Twins approach to estimate individualised treatment effects and then cluster them. As proposed in the literature, we use a survival tree to estimate response under each treatment regime and then use a KMeans (MacQueen et al., 1967) on the estimated treatment effects, computed as the difference between survival estimates. Note that these methodologies assume treatment randomisation and the clustering is uninformed by the covariates. For details on training and hyperparameter optimisation, we refer to App. B.2.
@@ 4.3 Treatment effect recovery
Figure 4: Averaged negative log-likelihood across 5-fold cross-validation given the number of subgroups $K$ under the ”Observational” treatment assignment with the shaded area representing 95% CI. The log-likelihood presents an elbow around the underlying number of subgroups.
@@
For all methodologies, we must choose the number of subgroups a priori. An important question is, therefore, whether we can identify the true number of subgroups with our model. Figure 4 presents the average negative log-likelihood obtained by cross-validation for models with different numbers of subgroups $K$. The dotted lines represent the elbow heuristic (Thorndike, 1953), which identifies a change point in the explained variability of a clustering strategy, here considering the log-likelihood. Using this heuristic, the optimal choice for $K$ is $3$, which aligns with the underlying generative process. App. B.3 shows that when the number of clusters is misspecified, the estimated treatment effects are unstable, the variance associated with estimated groups is hence an additional tool to select the number of clusters. This data-driven choice of the number of subgroups $K$ is a crucial strength of our method, compared to classical two-stage analyses, which separate clustering from treatment effect estimates. In these methods, survival outcomes cannot directly guide the choice of $K$, whereas it is directly reflected in the proposed methodology’s performance.
@@
Next, we present the quantitative results of our method against the benchmark methods. Table 1 summarises the performance of the different methodologies under the two studied scenarios. Recall that $K$ denotes the number of treatment response subgroups, while in CMHE, the additional parameter $L$ describes the number of survival clusters.
@@
Model Rand-Index $\text{IAE}_{k}(t_{max})$ $k=1$ $k=2$ $k=3$ Randomised CSC 0.807 (0.055) 0.037 (0.008) 0.023 (0.008) 0.015 (0.003) CSC Unadjusted 0.802 (0.050) 0.039 (0.011) 0.021 (0.009) 0.014 (0.003) CMHE ($K=3$) 0.392 (0.034) 0.164 (0.009) 0.077 (0.004) 0.070 (0.005) CMHE ($L=3$) 0.255 (0.159) - 0.084 (0.005) - CMHE ($K=L$) 0.345 (0.104) 0.243 (0.042) - - KMeans + TE 0.000 (0.002) 0.210 (0.011) 0.091 (0.007) 0.142 (0.016) Virtual Twins 0.578 (0.081) 0.022 (0.006) 0.032 (0.011) 0.025 (0.010) Observational CSC 0.797 (0.043) 0.042 (0.011) 0.051 (0.026) 0.022 (0.004) CSC Unadjusted 0.742 (0.073) 0.047 (0.019) 0.030 (0.029) 0.020 (0.003) CMHE ($K=3$) 0.385 (0.022) 0.169 (0.012) 0.078 (0.005) 0.075 (0.008) CMHE ($L=3$) 0.190 (0.127) 0.192 (0.010) - 0.140 (0.009) CMHE ($K=L$) 0.454 (0.068) 0.210 (0.013) 0.095 (0.020) 0.188 (0.014) KMeans + TE 0.001 (0.004) 0.192 (0.016) 0.090 (0.019) 0.147 (0.016) Virtual Twins 0.438 (0.139) 0.050 (0.039) 0.034 (0.033) 0.051 (0.033)
@@
Table 1: 5-fold cross-validated performance averaged (with standard deviation in parenthesis) under the Randomised and Observational treatment simulation. Best performance per column and simulation scenario are marked in bold, second best in italic. ’-’ describes when the methodology diverges. Our proposed CSC method and its unadjusted variant best recover the underlying treatment responses, with the adjusted approach presenting the best performance in the observational setting.
@@
CSC outperforms all CMHE alternatives, the current state-of-the-art method. CMHE’s parametrisation and assumptions explain this difference. Critically, CMHE assumes (i) proportional hazards and (ii) a linear impact of treatment on log hazards. Neither of these two assumptions is likely to hold in real-world settings as mimicked by our synthetic data. In contrast, our approach does not constrain the treatment effect due to its flexible modelling of the survival function under both treatment regimes. Increasing the number of subgroups, as shown in $K=L$, improves the performance of CMHE in terms of clustering quality (Rand-Index) and the recovery of the underlying treatment effect (lower IAE), but still is inferior compared to our proposed methods. These experiments highlight the advantage of the proposed CSC method in uncovering subgroups of treatment responses due to the flexibility in modelling complex survival distributions under both treatment regimes without proportionality assumption.
@@
CSC presents the best performance in identifying the underlying subgroup. Our proposed CSC method outperforms all approaches on the Rand-Index. In particular, the two-step approaches do not identify the underlying subgroups when covariates and outcomes subgroups are unaligned. KMeans + TE identify subgroups that are independent of the outcome of interest, as shown by the Rand-Index. Consequently, the identified clusters differ in their treatment responses, as shown by the large IAE. The Virtual Twins approach presents a better capacity to identify the clusters of interest with better Rand-Index and IAE. However, the two-step approach hurts subgroup identification due to the disconnect between the clustering and input covariates. This disconnect leads to a significantly lower Rand-Index in comparison to CSC despite accurate treatment effect estimates.
@@
Treatment assignment correction improves subgroup identification in observational settings. The Observational simulation demonstrates the importance of correcting the likelihood under treatment non-randomisation. The two CSC alternatives present comparable performance in the randomised setting, as theoretically expected, due to a constant $w_{i}$ in this context. However, CSC better recovers the different groups as shown by the Rand-Index. Finally, Appendices B.5.1 to B.5.4 further validate the robustness of our method in settings with unequally distributed cluster sizes, population size, increased number of underlying clusters, different treatment rates and covariate structures.
@@
Using the Surveillance, Epidemiology, and End Results program(Seer) [^7] Available at https://seer.cancer.gov/, App. C investigates a real-world setting: we analyse which subgroups of breast cancer patients most benefit from adjuvant radiotherapy, an active area of research. Our analysis highlights a subgroup characterised by a higher number of distant lymph nodes that benefit from treatment. Despite the inherent limitations of the dataset, our exploratory method identifies a subgroup that alternative methods miss, potentially informing future clinical trials to validate or challenge this hypothesis. Critically, this demonstrates that our work offers a practical methodology for exploratory subgroup analysis of medical data to identify treatment effect heterogeneity in observational studies.